1. Question: By partial fraction method, find the inverse z-transform of X(z): X(a)=7-4. and x[n] is stable. signals and systems
Answer #1

Xcz) = 3 (82) 822-22-) zu Z X(z) - 8 z 27-1 (x ) ३५ 87 27- XC) xezer 1 (2-1) XCX) 2 + 2] 1 - A(az-1) + B(47+1) 2A + 4B = 0 -AX(Z) 24 47+1 27-1 X(Z) -16 4 7+1 . 27-1 XC) - / 10 X(Z)= xce)= 4 Z u Z - Z- + Inverse - transform Xeco)4 (+)’und - 4 (44)”un)